# LTM: Scalable and Black-box Similarity-based Test Suite Minimization based on Language Models

### v3

@@ 4.2.2 Dataset
TABLE I: Summary of the $16$ Java projects
@@ 4.3 Results
4.3.1 A preliminary analysis of using full test code versus preprocessed test code
One important question is whether the source code of test cases requires preprocessing, such as removing test case names, logging statements, comments, and test assertions. Though such preprocessing was performed by ATM [8], this information may still convey relevant information about test case similarity, thus retaining them can improve fault detection capability. This information can also help language models better learn about semantic and contextual information about test cases and thus generates more informative test method embeddings. For example, code comment ‘Tests the equals method’ describes the functionality of the test code, thus providing potentially important semantic information [12]. Therefore, we conducted a preliminary study to compare the FDR results using full code (i.e., the source code of test cases with no preprocessing) and preprocessed code (as done by ATM) on a sample of $10$ projects from the Defects4J dataset. We observed that LTM achieved higher FDR results using full code than using prepossessed code ($+0.06$ on average) across all configurations, thus making it a better choice.
Note that we did not consider using the full code for ATM. Using unprocessed code would make the preparation phase of ATM much more time- and resource-consuming. This is due to the fact that (1) the xml files used to store Abstract Syntax Trees (ASTs) of test cases require substantial storage space and (2) it would take longer time to traverse the ASTs of test cases, thus yielding longer preparation time for transforming the test case code into ASTs and calculating similarity values. Given that, for ATM, preparation time can take up to 41.2% of the total minimization time, using the full code for ATM would make matters much worse.
@@ 4.3.1 A preliminary analysis of using full test code versus preprocessed test code
As discussed in Section 3.3, we optimized GA to accelerate the search process. Our results revealed that, compared to MT before optimization, the search time of ATM/Combined and ATM/TreeEditDistance is $224.66$ and $179.30$ times faster per project version, respectively. For LTM with UniXcoder/Cosine, the search time is $272.90$ times faster per project version, a significant improvement in scalability. Note that the following comparisons between LTM and ATM are based on applying this optimized GA to both approaches.
@@ 4.3.3 RQ1 results
We observed that, for each language model, there seems to be a significant difference in FDR between using Cosine Similarity and Euclidean Distance, which is further confirmed by Fisher’s exact test with $\alpha=0.01$. This suggests that FDR results are influenced by the combination of language models and distance functions. The embeddings generated by UniXcoder and GraphCodeBERT are enriched with detailed information about the nodes and edges within test code ASTs. This could be the reason that Cosine similarity, as a directional similarity, is more informative in distinguishing such test code embeddings than Euclidean Distance, a magnitude similarity. For CodeLlama-7b, the embedding dimensionality is $4,096$, compared to $768$ for other language models, which can explain its better performance with Cosine Similarity over Euclidean Distance, since the latter is less effective in high-dimensional spaces [46].
In certain projects ($Cli$, $Codec$, $Csv$, and $Time$), UniXcoder/Cosine yielded a $5\%$ to $10\%$ lower FDR than other configurations. However, Fisher’s Exact Test revealed that those differences are not significant. Nevertheless, we provide some tentative explanations regarding these projects: (1) A larger training data set covering a greater variety of programming languages (StarEncoder) and a substantially greater number of parameters (CodeLlama) might be beneficial for understanding test code from these projects, and (2) using ASTs as input for pre-training might introduce unnecessary details into test code embeddings, if test cases in these projects are similar in terms of code structure but rather differ regarding low-level details (e.g., complex string manipulation or domain-specific character strings), thus making UniXcoder less effective in understanding their test code.
In addition, we observed that UniXcoder/Cosine achieved higher average FDR than CodeLlama-7b/Cosine (mean = $0.80$, median = $0.81$) and CodeLlama-7b/Euclidean (mean = $0.76$, median = $0.76$). This suggests that language models with fewer parameters ($125Million$ for UniXcoder) can outperform much larger language models ($7Billion$ for CodeLlama-7b) for our minimization problem. This can be attributed to the fact that CodeLlama-7b, like most large language models, is a decoder-only model that was pre-trained for code generation tasks, while UniXcoder employs code understanding tasks that are more effective in analyzing and comprehending the contextual semantic information from the source code [47].
@@
 Project Approach CodeBERT GraphCodeBERT UniXcoder StarEncoder CodeLlama-7b ATM ATM
Cosine Euclidean Cosine Euclidean Cosine Euclidean Cosine Euclidean Cosine Euclidean Combined Tree Edit Distance
Chart 0.82 0.84 0.79 0.75 0.82 0.69 0.71 0.79 0.72 0.73 0.88 0.89
Cli 0.81 0.81 0.83 0.86 0.82 0.81 0.83 0.86 0.87 0.86 0.87 0.87
Codec 0.89 0.90 0.83 0.88 0.80 0.81 0.82 0.83 0.87 0.78 0.77 0.80
Collections 0.70 0.82 0.95 0.75 0.95 0.70 0.70 0.75 0.78 0.70 0.98 0.92
Compress 0.85 0.86 0.87 0.87 0.86 0.81 0.78 0.84 0.85 0.79 0.88 0.89
Csv 0.86 0.87 0.88 0.91 0.85 0.79 0.84 0.86 0.93 0.88 0.92 0.91
Gson 0.70 0.71 0.72 0.75 0.81 0.75 0.76 0.76 0.81 0.77 0.76 0.77
JacksonCore 0.70 0.73 0.73 0.73 0.75 0.67 0.74 0.74 0.75 0.75 0.69 0.72
JacksonDatabind 0.66 0.69 0.68 0.67 0.76 0.65 0.65 0.69 0.67 0.65 0.69 0.70
JacksonXml 0.77 0.85 0.83 0.73 0.83 0.82 0.67 0.65 0.83 0.85 0.60 0.73
Jsoup 0.75 0.74 0.77 0.76 0.80 0.76 0.73 0.75 0.77 0.75 0.70 0.69
JxPath 0.88 0.90 0.92 0.86 0.94 0.83 0.72 0.77 0.78 0.75 0.79 0.79
Lang 0.81 0.85 0.86 0.75 0.90 0.71 0.77 0.85 0.83 0.71 0.78 0.83
Math 0.74 0.78 0.75 0.74 0.78 0.65 0.71 0.78 0.75 0.67 0.81 0.82
Mockito 0.80 0.83 0.84 0.81 0.91 0.79 0.77 0.83 0.80 0.77 0.77 0.79
Time 0.83 0.88 0.85 0.82 0.82 0.77 0.82 0.79 0.81 0.76 0.89 0.88
@@
Min 0.66 0.69 0.68 0.67 0.75 0.65 0.65 0.65 0.67 0.65 0.60 0.69
25% Quantile 0.73 0.77 0.77 0.75 0.80 0.70 0.71 0.75 0.77 0.73 0.75 0.76
Mean 0.79 0.82 0.82 0.79 0.84 0.75 0.75 0.78 0.80 0.76 0.80 0.81
Median 0.81 0.84 0.83 0.76 0.82 0.77 0.75 0.79 0.81 0.76 0.79 0.81
75% Quantile 0.84 0.86 0.86 0.86 0.87 0.81 0.79 0.83 0.84 0.78 0.88 0.88
Max 0.89 0.90 0.95 0.91 0.95 0.83 0.84 0.86 0.93 0.88 0.98 0.92
@@
 Project Approach CodeBERT GraphCodeBERT UniXcoder StarEncoder CodeLlama-7b ATM ATM
Cosine Euclidean Cosine Euclidean Cosine Euclidean Cosine Euclidean Cosine Euclidean Combined Tree Edit Distance
Chart 0.92 1.51 0.97 1.20 1.33 0.99 2.24 2.68 21.02 20.92 7.86 94.22
Cli 0.11 0.16 0.14 0.15 0.17 0.14 0.16 0.22 4.39 4.38 0.23 1.20
Codec 0.15 0.21 0.18 0.17 0.22 0.18 0.25 0.33 5.75 5.65 0.42 3.02
Collections 0.37 0.51 0.43 0.50 0.58 0.38 0.79 1.11 8.96 8.80 2.16 14.89
Compress 0.16 0.20 0.19 0.18 0.22 0.17 0.25 0.34 9.26 9.18 0.55 6.55
Csv 0.10 0.13 0.12 0.11 0.15 0.11 0.14 0.16 4.19 4.15 0.16 0.35
Gson 0.31 0.39 0.31 0.33 0.47 0.32 0.68 0.79 8.31 8.26 1.48 6.09
JacksonCore 0.14 0.16 0.16 0.15 0.19 0.15 0.22 0.25 11.10 11.06 0.35 2.13
JacksonDatabind 0.92 1.40 0.91 1.35 1.43 1.33 2.15 2.46 34.67 35.41 4.90 22.69
JacksonXml 0.11 0.11 0.11 0.10 0.13 0.11 0.12 0.13 6.86 6.84 0.13 0.34
Jsoup 0.17 0.17 0.17 0.17 0.22 0.18 0.29 0.29 20.92 20.91 0.55 1.63
JxPath 0.11 0.15 0.13 0.13 0.17 0.13 0.17 0.21 4.98 4.95 0.22 0.73
Lang 0.91 1.47 0.91 1.08 1.38 0.97 2.35 2.82 15.09 14.98 5.44 33.47
Math 1.33 2.33 1.31 2.06 2.12 1.98 3.41 4.22 21.12 23.09 12.05 215.43
Mockito 0.40 0.50 0.42 0.40 0.59 0.41 0.97 1.12 17.40 17.93 2.32 9.34
Time 3.49 5.35 3.43 3.91 3.73 3.44 9.06 10.45 41.75 39.60 26.13 163.18
@@
Min 0.10 0.11 0.11 0.10 0.13 0.11 0.12 0.13 4.19 4.15 0.13 0.34
25% Quantile 0.13 0.16 0.16 0.15 0.19 0.15 0.21 0.24 6.58 6.54 0.32 1.52
Mean 0.61 0.92 0.68 0.75 0.82 0.69 1.45 1.72 14.74 14.72 4.06 35.95
Median 0.24 0.30 0.25 0.26 0.35 0.25 0.49 0.57 10.18 10.12 1.02 6.32
75% Quantile 0.91 1.42 0.91 1.11 1.34 0.98 2.17 2.52 20.95 20.91 5.04 25.39
Max 3.49 5.35 3.43 3.91 3.73 3.44 9.06 10.45 41.75 39.60 26.13 215.43
@@
 Project Approach CodeBERT GraphCodeBERT UniXcoder StarEncoder CodeLlama-7b ATM ATM
Cosine Euclidean Cosine Euclidean Cosine Euclidean Cosine Euclidean Cosine Euclidean Combined Tree Edit Distance
Chart 55.45% 57.35% 58.78% 59.81% 58.03% 56.11% 58.47% 63.08% 56.64% 59.14% 59.77% 57.53%
Cli 29.90% 27.90% 30.54% 29.75% 28.32% 28.17% 29.52% 27.52% 28.25% 28.46% 28.07% 27.34%
Codec 44.11% 40.68% 47.59% 49.93% 55.58% 50.62% 51.05% 48.38% 45.27% 48.83% 34.62% 45.94%
Collections 57.40% 55.05% 56.21% 56.10% 55.41% 59.05% 59.41% 57.36% 55.01% 56.78% 53.57% 57.65%
Compress 37.74% 37.98% 36.56% 38.10% 39.10% 40.70% 37.87% 35.26% 36.45% 40.11% 33.85% 32.59%
Csv 18.66% 18.78% 19.90% 20.07% 18.13% 10.11% 16.40% 16.15% 12.82% 9.37% 20.45% 19.89%
Gson 41.81% 40.60% 42.16% 42.01% 41.40% 44.63% 41.85% 40.10% 41.69% 43.63% 40.35% 38.17%
JacksonCore 47.45% 48.44% 44.37% 46.18% 47.10% 48.00% 46.02% 46.42% 46.95% 46.98% 43.59% 46.71%
JacksonDatabind 42.54% 41.28% 41.68% 41.63% 38.32% 42.37% 41.71% 39.56% 41.29% 41.75% 50.61% 47.92%
JacksonXml 42.81% 42.05% 42.63% 44.12% 45.80% 46.32% 45.70% 49.52% 44.03% 43.60% 39.17% 39.34%
Jsoup 42.96% 42.03% 41.33% 43.80% 33.90% 41.65% 38.60% 35.61% 35.12% 37.55% 40.39% 38.20%
JxPath 43.37% 42.82% 41.00% 41.90% 40.83% 42.70% 42.46% 41.84% 44.12% 44.16% 44.40% 44.46%
Lang 38.36% 36.49% 40.14% 41.24% 34.05% 44.76% 41.04% 35.37% 42.83% 44.97% 37.71% 38.06%
Math 50.24% 49.43% 50.26% 49.47% 49.67% 49.67% 50.00% 49.72% 50.90% 51.34% 47.93% 49.10%
Mockito 27.65% 29.87% 34.33% 38.38% 33.97% 39.93% 44.82% 41.80% 32.13% 39.77% 24.31% 26.13%
Time 48.12% 47.57% 48.33% 48.49% 47.89% 49.06% 47.35% 46.94% 46.84% 48.38% 45.93% 47.35%
@@
Min 18.66% 18.78% 19.90% 20.07% 18.13% 10.11% 16.40% 16.15% 12.82% 9.37% 20.45% 19.89%
25% Quantile 38.21% 37.61% 39.25% 40.53% 34.03% 41.41% 40.43% 35.55% 36.12% 40.03% 34.43% 36.69%
Mean 41.79% 41.15% 42.24% 43.19% 41.72% 43.37% 43.27% 42.16% 41.27% 42.80% 40.30% 41.02%
Median 42.89% 41.66% 41.92% 42.91% 41.12% 44.70% 43.64% 41.82% 43.43% 43.90% 40.37% 41.90%
75% Quantile 47.62% 47.79% 47.78% 48.74% 48.34% 49.21% 48.01% 48.67% 46.87% 48.49% 46.43% 47.49%
Max 57.40% 57.35% 58.78% 59.81% 58.03% 59.05% 59.41% 63.08% 56.64% 59.14% 59.77% 57.65%
@@
Search Time. The average search time of LTM ranges from $0.46$ to $1.90$ min per project version, across configurations, with a maximum of $24.49$ min. LTM using CodeBERT with Cosine similarity was the fastest in terms of search time, ranging from $0.04$ to $3.86$ min, across project versions. Not only LTM using UniXcoder/Cosine took a shorter search time ($0.78$ min) than the average search time across configurations ($1.03$ min), it also achieved the highest FDR among all configurations, thus making it the best option for LTM.
Results for the Closure project. We further evaluated LTM on the $Closure$ project, which has $174$ versions with an average of $7,308$ test cases per version. Though this project is part of the Defects4J dataset, it was previously excluded from the ATM evaluation due its scalability limitations [8] and the large size of our experiments. However, evaluating the effectiveness and efficiency of LTM on this project, considering its larger scale and test suite, can be informative. Table V presents the results regarding FDR, MT (in min), and TSR (in percentage), respectively. We can see that UniXcoder/Cosine achieves once again the highest FDR of $0.79$ and an MT of only $17.90~{}min$, while saving $52.55\%$ testing time, which indicates that LTM can be effective and scalable for larger projects and test suites.
RQ1 summary. LTM achieved high FDR results ($0.79$, on average, across configurations) for a $50\%$ minimization budget. UniXcoder/Cosine is the best LTM configuration when considering both effectiveness ($0.84$ FDR on average) and efficiency ($0.82$ min on average), with an average TSR of $41.72\%$

### v4

@@ 4.2.2 Dataset
TABLE I: Summary of the $17$ Java projects
@@ 4.3 Results
4.3.1 Using full test code versus preprocessed test code in LTM
One important question is whether the source code of test cases requires preprocessing, such as removing test case names, logging statements, comments, and test assertions. Though such preprocessing was performed by ATM [8], due to necessity (Section 4.2.1), this information may still convey relevant information about test case similarity, thus retaining them can potentially improve fault detection capability. This information can also help language models better learn about semantic and contextual information about test cases and thus generate more informative test method embeddings. For example, code comment ‘Tests the equals method’ describes the functionality of the test code, thus providing potentially important semantic information [12]. Therefore, we conducted a preliminary study to compare the FDR results of LTM using full code (i.e., the source code of test cases without preprocessing) and preprocessed code on a sample of $10$ smaller projects from the Defects4J dataset. We observed that LTM achieved higher FDR results using full code than using prepossessed code ($+0.02$ on average) across all configurations, thus making the former a better choice for our experiments.
@@ 4.3.1 Using full test code versus preprocessed test code in LTM
As discussed in Section 3.3, we optimized GA to accelerate the search process. Our results revealed that, compared to MT before optimization, the search time of ATM/Combined and ATM/TreeEditDistance is $224.66$ and $203.02$ times faster per project version, respectively. For LTM, the search time is $190.48$ times faster per project version, across all configurations, which is a significant improvement in scalability. Note that the following comparisons between LTM and ATM are based on applying this optimized GA to both approaches.
@@ 4.3.3 RQ1 results
We observed that, for each language model, there seems to be a significant difference in FDR between using Cosine Similarity and Euclidean Distance, which is further confirmed by Fisher’s exact test with $\alpha=0.01$. This suggests that FDR results are influenced by the combination of language models and distance functions. The embeddings generated by UniXcoder and GraphCodeBERT are enriched with detailed information about the nodes and edges within test code ASTs. This could be the reason that Cosine Similarity, as a directional similarity, is more informative in distinguishing such test method embeddings than Euclidean Distance, a magnitude similarity. For CodeLlama-7b, the embedding dimensionality is $4,096$, compared to $768$ for other language models, which can explain its better performance with Cosine Similarity over Euclidean Distance, since the latter is less effective in high-dimensional spaces [46].
In certain projects ($Cli$, $Codec$, $Csv$, and $Time$), UniXcoder/Cosine yielded a $5\%$ to $10\%$ lower FDR than other configurations. However, Fisher’s Exact Test revealed that those differences are not significant. Nevertheless, we provide some tentative explanations regarding these projects: (1) A larger training data set covering a greater variety of programming languages (StarEncoder) and a substantially greater number of parameters (CodeLlama) might be beneficial for understanding test code from these projects, and (2) using ASTs as input for pre-training might introduce unnecessary details into test method embeddings, if test cases in these projects are similar in terms of code structure but rather differ regarding low-level details (e.g., complex string manipulation or domain-specific character strings), thus making UniXcoder less effective in understanding their test code.
In addition, we observed that UniXcoder/Cosine achieved a higher average FDR than CodeLlama-7b/Cosine (mean = $0.80$, median = $0.81$) and CodeLlama-7b/Euclidean (mean = $0.76$, median = $0.76$). This suggests that language models with fewer parameters ($125Million$ for UniXcoder) can outperform much larger language models ($7Billion$ for CodeLlama-7b) for our minimization problem. This can be attributed to the fact that CodeLlama-7b, like most large language models, is a decoder-only model that was pre-trained for code generation tasks, while UniXcoder employs code understanding tasks that are more effective in analyzing and comprehending the contextual semantic information from the source code [47].
@@
 Project Approach CodeBERT GraphCodeBERT UniXcoder StarEncoder CodeLlama-7b UniXcoder-p ATM ATM
Cosine Euclidean Cosine Euclidean Cosine Euclidean Cosine Euclidean Cosine Euclidean Cosine Combined Tree Edit Distance
Chart 0.82 0.84 0.79 0.75 0.82 0.69 0.71 0.79 0.72 0.73 0.77 0.88 0.89
Cli 0.81 0.81 0.83 0.86 0.82 0.81 0.83 0.86 0.87 0.86 0.80 0.87 0.87
Codec 0.89 0.90 0.83 0.88 0.80 0.81 0.82 0.83 0.87 0.78 0.78 0.77 0.80
Collections 0.70 0.82 0.95 0.75 0.95 0.70 0.70 0.75 0.78 0.70 1.00 0.98 0.92
Compress 0.85 0.86 0.87 0.87 0.86 0.81 0.78 0.84 0.85 0.79 0.89 0.88 0.89
Csv 0.86 0.87 0.88 0.91 0.85 0.79 0.84 0.86 0.93 0.88 0.88 0.92 0.91
Gson 0.70 0.71 0.72 0.75 0.81 0.75 0.76 0.76 0.81 0.77 0.76 0.76 0.77
JacksonCore 0.70 0.73 0.73 0.73 0.75 0.67 0.74 0.74 0.75 0.75 0.78 0.69 0.72
JacksonDatabind 0.66 0.69 0.68 0.67 0.76 0.65 0.65 0.69 0.67 0.65 0.74 0.69 0.70
JacksonXml 0.77 0.85 0.83 0.73 0.83 0.82 0.67 0.65 0.83 0.85 0.87 0.60 0.73
Jsoup 0.75 0.74 0.77 0.76 0.80 0.76 0.73 0.75 0.77 0.75 0.76 0.70 0.69
JxPath 0.88 0.90 0.92 0.86 0.94 0.83 0.72 0.77 0.78 0.75 0.94 0.79 0.79
Lang 0.81 0.85 0.86 0.75 0.90 0.71 0.77 0.85 0.83 0.71 0.84 0.78 0.83
Math 0.74 0.78 0.75 0.74 0.78 0.65 0.71 0.78 0.75 0.67 0.77 0.81 0.82
Mockito 0.80 0.83 0.84 0.81 0.91 0.79 0.77 0.83 0.80 0.77 0.78 0.77 0.79
Time 0.83 0.88 0.85 0.82 0.82 0.77 0.82 0.79 0.81 0.76 0.86 0.89 0.88
@@
Min 0.66 0.69 0.68 0.67 0.75 0.65 0.65 0.65 0.67 0.65 0.74 0.60 0.69
25% Quantile 0.73 0.77 0.77 0.75 0.80 0.70 0.71 0.75 0.77 0.73 0.77 0.75 0.76
Mean 0.79 0.82 0.82 0.79 0.84 0.75 0.75 0.78 0.80 0.76 0.83 0.80 0.81
Median 0.81 0.84 0.83 0.76 0.82 0.77 0.75 0.79 0.81 0.76 0.79 0.79 0.81
75% Quantile 0.84 0.86 0.86 0.86 0.87 0.81 0.79 0.83 0.84 0.78 0.87 0.88 0.88
Max 0.89 0.90 0.95 0.91 0.95 0.83 0.84 0.86 0.93 0.88 1.00 0.98 0.92
@@
 Project Approach CodeBERT GraphCodeBERT UniXcoder StarEncoder CodeLlama-7b UniXcoder-p ATM ATM
Cosine Euclidean Cosine Euclidean Cosine Euclidean Cosine Euclidean Cosine Euclidean Cosine Combined Tree Edit Distance
Chart 0.92 1.51 0.97 1.20 1.33 0.99 2.24 2.68 21.02 20.92 2.18 7.86 94.22
Cli 0.11 0.16 0.14 0.15 0.17 0.14 0.16 0.22 4.39 4.38 0.22 0.23 1.20
Codec 0.15 0.21 0.18 0.17 0.22 0.18 0.25 0.33 5.75 5.65 0.31 0.42 3.02
Collections 0.37 0.51 0.43 0.50 0.58 0.38 0.79 1.11 8.96 8.80 0.87 2.16 14.89
Compress 0.16 0.20 0.19 0.18 0.22 0.17 0.25 0.34 9.26 9.18 0.32 0.55 6.55
Csv 0.10 0.13 0.12 0.11 0.15 0.11 0.14 0.16 4.19 4.15 0.19 0.16 0.35
Gson 0.31 0.39 0.31 0.33 0.47 0.32 0.68 0.79 8.31 8.26 0.77 1.48 6.09
JacksonCore 0.14 0.16 0.16 0.15 0.19 0.15 0.22 0.25 11.10 11.06 0.27 0.35 2.13
JacksonDatabind 0.92 1.40 0.91 1.35 1.43 1.33 2.15 2.46 34.67 35.41 2.50 4.90 22.69
JacksonXml 0.11 0.11 0.11 0.10 0.13 0.11 0.12 0.13 6.86 6.84 0.18 0.13 0.34
Jsoup 0.17 0.17 0.17 0.17 0.22 0.18 0.29 0.29 20.92 20.91 0.33 0.55 1.63
JxPath 0.11 0.15 0.13 0.13 0.17 0.13 0.17 0.21 4.98 4.95 0.23 0.22 0.73
Lang 0.91 1.47 0.91 1.08 1.38 0.97 2.35 2.82 15.09 14.98 2.53 5.44 33.47
Math 1.33 2.33 1.31 2.06 2.12 1.98 3.41 4.22 21.12 23.09 3.69 12.05 215.43
Mockito 0.40 0.50 0.42 0.40 0.59 0.41 0.97 1.12 17.40 17.93 1.15 2.32 9.34
Time 3.49 5.35 3.43 3.91 3.73 3.44 9.06 10.45 41.75 39.60 7.61 26.13 163.18
@@
Min 0.10 0.11 0.11 0.10 0.13 0.11 0.12 0.13 4.19 4.15 0.18 0.13 0.34
25% Quantile 0.13 0.16 0.16 0.15 0.19 0.15 0.21 0.24 6.58 6.54 0.26 0.32 1.52
Mean 0.61 0.92 0.68 0.75 0.82 0.69 1.45 1.72 14.74 14.72 1.46 4.06 35.95
Median 0.24 0.30 0.25 0.26 0.35 0.25 0.49 0.57 10.18 10.12 0.55 1.02 6.32
75% Quantile 0.91 1.42 0.91 1.11 1.34 0.98 2.17 2.52 20.95 20.91 2.26 5.04 25.39
Max 3.49 5.35 3.43 3.91 3.73 3.44 9.06 10.45 41.75 39.60 7.61 26.13 215.43
@@
 Project Approach CodeBERT GraphCodeBERT UniXcoder StarEncoder CodeLlama-7b UniXcoder-p ATM ATM
Cosine Euclidean Cosine Euclidean Cosine Euclidean Cosine Euclidean Cosine Euclidean Cosine Combined Tree Edit Distance
Chart 55.45% 57.35% 58.78% 59.81% 58.03% 56.11% 58.47% 63.08% 56.64% 59.14% 55.27% 59.77% 57.53%
Cli 29.90% 27.90% 30.54% 29.75% 28.32% 28.17% 29.52% 27.52% 28.25% 28.46% 29.95% 28.07% 27.34%
Codec 44.11% 40.68% 47.59% 49.93% 55.58% 50.62% 51.05% 48.38% 45.27% 48.83% 45.84% 34.62% 45.94%
Collections 57.40% 55.05% 56.21% 56.10% 55.41% 59.05% 59.41% 57.36% 55.01% 56.78% 56.37% 53.57% 57.65%
Compress 37.74% 37.98% 36.56% 38.10% 39.10% 40.70% 37.87% 35.26% 36.45% 40.11% 36.33% 33.85% 32.59%
Csv 18.66% 18.78% 19.90% 20.07% 18.13% 10.11% 16.40% 16.15% 12.82% 9.37% 6.92% 20.45% 19.89%
Gson 41.81% 40.60% 42.16% 42.01% 41.40% 44.63% 41.85% 40.10% 41.69% 43.63% 39.51% 40.35% 38.17%
JacksonCore 47.45% 48.44% 44.37% 46.18% 47.10% 48.00% 46.02% 46.42% 46.95% 46.98% 45.22% 43.59% 46.71%
JacksonDatabind 42.54% 41.28% 41.68% 41.63% 38.32% 42.37% 41.71% 39.56% 41.29% 41.75% 39.98% 50.61% 47.92%
JacksonXml 42.81% 42.05% 42.63% 44.12% 45.80% 46.32% 45.70% 49.52% 44.03% 43.60% 44.48% 39.17% 39.34%
Jsoup 42.96% 42.03% 41.33% 43.80% 33.90% 41.65% 38.60% 35.61% 35.12% 37.55% 32.55% 40.39% 38.20%
JxPath 43.37% 42.82% 41.00% 41.90% 40.83% 42.70% 42.46% 41.84% 44.12% 44.16% 41.04% 44.40% 44.46%
Lang 38.36% 36.49% 40.14% 41.24% 34.05% 44.76% 41.04% 35.37% 42.83% 44.97% 32.19% 37.71% 38.06%
Math 50.24% 49.43% 50.26% 49.47% 49.67% 49.67% 50.00% 49.72% 50.90% 51.34% 45.44% 47.93% 49.10%
Mockito 27.65% 29.87% 34.33% 38.38% 33.97% 39.93% 44.82% 41.80% 32.13% 39.77% 24.12% 24.31% 26.13%
Time 48.12% 47.57% 48.33% 48.49% 47.89% 49.06% 47.35% 46.94% 46.84% 48.38% 46.17% 45.93% 47.35%
@@
Min 18.66% 18.78% 19.90% 20.07% 18.13% 10.11% 16.40% 16.15% 12.82% 9.37% 6.92% 20.45% 19.89%
25% Quantile 38.21% 37.61% 39.25% 40.53% 34.03% 41.41% 40.43% 35.55% 36.12% 40.03% 32.46% 34.43% 36.69%
Mean 41.79% 41.15% 42.24% 43.19% 41.72% 43.37% 43.27% 42.16% 41.27% 42.80% 38.84% 40.30% 41.02%
Median 42.89% 41.66% 41.92% 42.91% 41.12% 44.70% 43.64% 41.82% 43.43% 43.90% 40.51% 40.37% 41.90%
75% Quantile 47.62% 47.79% 47.78% 48.74% 48.34% 49.21% 48.01% 48.67% 46.87% 48.49% 45.54% 46.43% 47.49%
Max 57.40% 57.35% 58.78% 59.81% 58.03% 59.05% 59.41% 63.08% 56.64% 59.14% 56.37% 59.77% 57.65%
@@
Search Time. The average search time of LTM ranges from $0.46$ to $1.90$ min per project version, across configurations. LTM using CodeBERT with Cosine similarity was the fastest in terms of search time, ranging from $0.04$ to $3.86$ min, across project versions. Not only LTM using UniXcoder/Cosine took a shorter search time ($0.78$ min) than the average search time across configurations ($1.03$ min), it also achieved the highest FDR among all configurations, thus making it the best option for LTM.
Results for the Closure project. We further evaluated LTM on the $Closure$ project, which has $174$ versions with an average of $7,308$ test cases per version. Though this project is part of the Defects4J dataset, it was previously excluded from the ATM evaluation due to its scalability limitations [8] and the large size of our experiments. However, evaluating the effectiveness and efficiency of LTM on this project, considering its larger scale and test suite, can be informative. Table V presents the results regarding FDR, MT (in min), and TSR (in percentage), respectively. We can see that UniXcoder/Cosine achieves once again the highest FDR of $0.79$ and an MT of only $17.80~{}min$, while saving $52.55\%$ testing time, which indicates that LTM can be effective and scalable for larger projects and test suites.
RQ1 summary. LTM achieved high FDR results ($0.79$, on average, across configurations) for a $50\%$ minimization budget. UniXcoder/Cosine is the best LTM configuration when considering both effectiveness ($0.84$ FDR on average) and efficiency ($0.82$ min on average), with an average TSR of $41.72\%$.